OpenStudy (anonymous):

What fraction is the same as 62/8? 8 1/4 7 3/8 7 5/8 Not Here 7 1/8

OpenStudy (saifoo.khan):

i would choose, not here.

OpenStudy (anonymous):

62/8 = 31/4 = 7 3/4, so "not here" is the answer